Student council consists of 3 seniors, 3 juniors, 5 sophomores and 3 freshman. A committee must be made up of two students from 9, 10 or 11th grade and another 2 students from 12th grade. How many different committees can be made? Explain. IT does not say whether seniors and serve on more than one committee.

Answer by stanbon(75887)

Student council consists of 3 seniors, 3 juniors, 5 sophomores and 3 freshman. A committee must be made up of two students from 9, 10 or 11th grade and another 2 students from 12th grade. How many different committees can be made? Explain. IT does not say whether seniors and serve on more than one committee.
====
Choose 2 of the 3 seniors in 3C2 = 3C1 = 3/1 = 3 ways
Choose 2 of the 11 underclassmen 11C2 = (11*10)/(1*2) = 55 ways
-------------------
Ans: 3C2*11C2 = 3*55 = 165
====
Cheers,
Stan H.
===========